One number is 4 less than another number. Find the numbers if three times the smaller number is 22 less than four times the larger number.
Answer by stanbon(75887)

One number is 4 less than another number. Find the numbers if three times the smaller number is 22 less than four times the larger number.
----
Equations:
a = b-4
3a = 4b-22
------
Substitute for "a" and solve for "b":
3(b-4) = 4b-22
--------
3b - 12 = 4b - 22
b = 10
----
a = b-4 = 6
===============
Cheers,
Stan H.
===================